Il Directions: Perform what is being asked Solve the problem by showing the step-by-step process.


Situation: Miko, a DOH COVID-19 tracker, found out that the number' fected persons of COVID-19 in seven


low nisk provinces of the country are 1,9, 2, 12, 8, 7, and 1 Supposed to provinces are drawn as sample,


1. What is the mean (), variance (o) and standard deation of the population (c)?


2 How many different samples of size n =6 can be drawn from the pop station?


List them with their corresponding mean


Construct the sampling distribution of the sample menn

1
Expert's answer
2022-03-14T17:10:08-0400

Solution

1.

For a population;

mean "\mu=\dfrac{\sum X_i}{N}"


Variance "\sigma^2 =\dfrac{\sum (X_i-\mu)^2}{N}"


Standard deviation "\sigma=\sqrt\dfrac{\sum(X_i-\mu)^2}{N}"


mean "\mu = \dfrac{1+9+2+12+8+7+1}{7}"


"\mu=5.71"


Variance


"\def\arraystretch{1.5}\begin{array}{c:c:c} X_i&X_i -\mu&(X_i-\mu)^2\\\hline1&-4.71&22.18\\\hdashline1&-4.71&22.18\\\hdashline2&-3.71&13.76\\\hdashline7&1.29&1.66\\\hdashline8&2.29&5.24\\\hdashline9&3.29&10.82\\\hdashline12&6.29&39.56\\\hline \sum&&115.40\\\hline\end{array}"


"\sigma^2=\dfrac{115.4}{7}=16.49"


Standard deviation

"\sigma=\sqrt\dfrac{115.4}{7} =4.06"


2.


Population size "N=7"

Sample size "n =6"

Possible number of samples

"=C_7^6= \binom{7}{6}=7" Samples


"\def\arraystretch{1.5}\begin{array}{c:c:c}no.& Sample & Sample mean\\\hline1&1,1,2,7,8,9&4.67\\\hdashline2&1,1,2,7,8,12&5.17\\\hdashline3&1,1,2,7,9,12&5.33\\\hdashline4&1,1,2,8,9,12&5.50\\\hdashline5&1,1,7,8,9,12&6.33\\\hdashline6&1,2,7,8,9,12&6.50\\\hdashline7&1,2,7,8,9,12&6.50\\\hline\end{array}"



Sampling distribution of the sample means


"\def\arraystretch{1.5}\begin{array}{c:c} Mean& Probability \\\hline4.67&\frac17\\\hdashline5.17&\frac17\\\hdashline5.33&\frac17\\\hdashline5.50&\frac17\\\hdashline6.33&\frac17\\\hdashline6.50&\frac27\\\hline\end{array}"